Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability and combinations

Expert replies
by venmic » Thu Apr 28, 2011 7:14 pm
can some one give me a complete explantion to his


There are 8 magazines lying on a table; 4 are fashion magazines and the other 4 are sports magazines. If 3 magazines are to be selected at random from the 8 magazines, what is the probability that at least one of the fashion magazines will be selected?

A. 1/2
B. 2/3
C. 32/35
D. 11/12
E. 13/14
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Thu Apr 28, 2011 8:05 pm
venmic wrote:can some one give me a complete explantion to his


There are 8 magazines lying on a table; 4 are fashion magazines and the other 4 are sports magazines. If 3 magazines are to be selected at random from the 8 magazines, what is the probability that at least one of the fashion magazines will be selected?

A. 1/2
B. 2/3
C. 32/35
D. 11/12
E. 13/14
Total no. of possible outcomes = 8C3 = 8 * 7 = 56 ways
If all 3 magazines are fashion magazines, then it can occur in 4C3 = 4 ways.
If 2 are fashion magazines, and 1 is sports magazine, then this can occur in 4C2 * 4C1 = 6 * 4 = 24 ways
If 1 is fashion magazine and 2 are sports magazines, then this can occur in 4C1 * 4C2 = 4 * 6 = 24 ways
So, no. of ways so that at least one of the fashion magazines is selected = 4 + 24 + 24 = 52 ways

Hence, required probability = 52/56 = 13/14

The correct answer is E.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by Stuart@KaplanGMAT » Fri Apr 29, 2011 12:18 am
venmic wrote:can some one give me a complete explantion to his


There are 8 magazines lying on a table; 4 are fashion magazines and the other 4 are sports magazines. If 3 magazines are to be selected at random from the 8 magazines, what is the probability that at least one of the fashion magazines will be selected?

A. 1/2
B. 2/3
C. 32/35
D. 11/12
E. 13/14
Hi,

on most complicated probability problems on the GMAT, the "one minus" approach is the quickest way to solve.

This approach is based on the equation:

Prob(what you want) + Prob(what you don't want) = 1

which rearranges to:

Prob(what you want) = 1 - Prob(what you don't want).

Applying that formula to this question, we see that we want at least 1 fashion magazine; so, what we don't want is 0 fashion magazines.

Prob(no fashion magazines) = Prob(3 sports magazines)

We can now solve using either combinations or straight probability.

Using combinations:

Prob = (# desired outcomes)/(total # of possibilities)
Prob(3 sports) = 4C3/8C3
Prob(3 sports) = (4/1)/(8*7*6/3*2)
Prob(3 sports) = 4/56 = 1/14

Prob(at least 1 fashion) = 1 - 1/14 = 13/14

Using straight probability:

Prob of multiple independent events all occurring = the product of the probabilities each event occurring

Prob(1st one sports)*Prob(2nd one sports)*Prob(3rd one sports)
= 4/8 * 3/7 * 2/6 = 4*3*2/8*7*6 = 4/8*7 = 4/56 = 1/14

Prob(at least 1 fashion) = 1 - 1/14 = 13/14
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by manpsingh87 » Fri Apr 29, 2011 1:09 am
venmic wrote:can some one give me a complete explantion to his


There are 8 magazines lying on a table; 4 are fashion magazines and the other 4 are sports magazines. If 3 magazines are to be selected at random from the 8 magazines, what is the probability that at least one of the fashion magazines will be selected?

A. 1/2
B. 2/3
C. 32/35
D. 11/12
E. 13/14
total no. of magazines=8;
out of 8 magazines any 3 can be selected in 8C3 ways=56;
probability that atleast one of the selected magazine is of fashion= 1- probability that none of the selected magazine is of fashion.

probability that none of the selected magazine is of fashion.= probability that all the selected magazine is of sports= 4C3;

hence required probability=1-4/56=13/14 hence E
O Excellence... my search for you is on... you can be far.. but not beyond my reach!
Join the discussion

by nitestr » Fri May 03, 2013 2:18 pm
Hi Stuart,

One quick question on this. By multiplying the Prob, is the order being counted as well?

Prob(1st one sports)*Prob(2nd one sports)*Prob(3rd one sports)
= 4/8 * 3/7 * 2/6 = 4*3*2/8*7*6 = 4/8*7 = 4/56 = 1/14


Thanks

ns
Join the discussion